Hi! I have a problem that says that the area of the trapezoid (which is 12, by the way...no units were indicated for some reason) is equal the area of the circle. What is the radius of the circle? My answer was sqrt(12/pi). However, I am very unsure of my answer. Am I right? Thanks for your time!
Answer by Alan3354(69443)

Area = 12.
----
Area of a circle = pi*r^2
pi*r^2 = 12
r^2 = 12/pi
r = sqrt(12/pi) --- in whatever units
============
You win.